### v1

@@ 2 $q$-Differential operators for $q$-spinor variables
The aim of this section is to define the $q$-differential operator for $q$- spinor variables. To begin, first we will mention an important rule of the $q$- spinor differential calculus, that is the q- spinor chain rule.
@@ 2.1 The $q$-spinor chain rule
Let us consider a $q$-spinor function on the form $\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu}))$. The $q$-spinor chain rule can be expressed as
$\frac{\partial^{q}\Psi}{\partial^{q}x_{\mu}}=\frac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial_{q}x_{\mu}}.$ (51)
@@
Consider the following derivative
$\frac{\partial^{q}\Psi}{\partial^{q}x_{\mu}}=\frac{\Psi((qu)^{\alpha}_{\dot{\beta}}(x_{\mu}))-q\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu}))}{(qx_{\mu})-q(x_{\mu})},$ (52)
multiplying by $\frac{(qu)^{\alpha}_{\dot{\beta}}(x_{\mu})-qu^{\alpha}_{\dot{\beta}}(x_{\mu})}{(qu)^{\alpha}_{\dot{\beta}}(x_{\mu})-qu^{\alpha}_{\dot{\beta}}(x_{\mu})}$, we can rewrite (52) as
$\frac{\partial^{q}\Psi}{\partial^{q}x_{\mu}}=\frac{\Psi((qu)^{\alpha}_{\dot{\beta}}(x_{\mu}))-q\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu}))}{(qu)^{\alpha}_{\dot{\beta}}(x_{\mu})-qu^{\alpha}_{\dot{\beta}}(x_{\mu})}\frac{(qu)^{\alpha}_{\dot{\beta}}(x_{\mu})-qu^{\alpha}_{\dot{\beta}}(x_{\mu})}{(qx_{\mu})-q(x_{\mu})},$ (53)
in virtue of (38) and denoting $\frac{(qu)^{\alpha}_{\dot{\beta}}(x_{\mu})-qu^{\alpha}_{\dot{\beta}}(x_{\mu})}{(qx_{\mu})-q(x_{\mu})}$ by $\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial^{q}x_{\mu}}$, finally we have
$\frac{\partial^{q}\Psi}{\partial^{q}x_{j}}=\frac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial_{q}x_{j}},$ (54)
@@
Now, we will define a new $q$-differential operator for $q$-spinor variables over an orthonormal basis of $\mathds{R}^{n}$, different to usual Dirac operator and Cauchy - Riemann operator mentioned in [3], [4], [6], and [7].
2.2 The new $q$-differential operator for $q$-spinor variables
The motivation comes from the construction of the any diiferential operator that satisfy the property $D_{q}^{2}=-\frac{\partial^{2}_{q}}{\partial_{q}x^{2}_{\mu}}-\frac{\partial^{2}_{q}}{\partial_{q}x^{2}_{\nu}}$ for all $\mu,\nu=1,2,...,n$ over an orthonormal basis of $\mathds{R}^{n}$.
Proposition 2.2.
Let $\left\{\mathrm{e}_{1},\mathrm{e}_{2},...,\mathrm{e}_{n}\right\}$ be an orthonormal basis of $\mathds{R}^{n}$. The $q$- Differential operator $D^{q}$
is defined over $\mathds{R}^{n}$ under the follows relations
$\mathrm{e}^{2}_{\mu}\frac{\partial_{q}^{2}}{\partial_{q}x^{2}_{\nu}}+\mathrm{e}^{2}_{\nu}\frac{\partial^{2}_{q}}{\partial_{q}x^{2}_{\mu}}$ $=\ -2\delta_{\mu\alpha}\frac{\partial^{2}_{q}}{\partial_{q}x_{\mu}\partial_{q}x_{\alpha}}-\frac{\partial^{2}_{q}}{\partial_{q}x^{2}_{\nu}},$ (56)
$\mathrm{e}_{\mu}\frac{\partial_{q}}{\partial_{q}x_{\nu}}\mathrm{e}_{\nu}\frac{\partial_{q}}{\partial_{q}x_{\mu}}+\mathrm{e}_{\nu}\frac{\partial_{q}}{\partial_{q}x_{\mu}}\mathrm{e}_{\mu}\frac{\partial_{q}}{\partial_{q}x_{\nu}}$ $=\ \delta_{\mu\alpha}\frac{\partial^{2}_{q}}{\partial_{q}x_{\mu}\partial_{q}x_{\alpha}},\quad\mu,\nu,\alpha=1,2,...n.$ (57)
Proof.
The proof is based on the following observation: the square of (55) is equivalent to $-\frac{\partial^{2}_{q}}{\partial_{q}x^{2}_{\mu}}-\frac{\partial^{2}_{q}}{\partial_{q}x^{2}_{\nu}}$. Therefore, we first compute $D^{2}_{q}$ resulting
$D^{2}_{q}$ $=\ \mathrm{e}^{2}_{\mu}\frac{\partial_{q}^{2}}{\partial_{q}x^{2}_{\nu}}+\mathrm{e}^{2}_{\nu}\frac{\partial^{2}_{q}}{\partial_{q}x^{2}_{\mu}}+\mathrm{e}_{\mu}\frac{\partial_{q}}{\partial_{q}x_{\nu}}\mathrm{e}_{\nu}\frac{\partial_{q}}{\partial_{q}x_{\mu}}+\mathrm{e}_{\nu}\frac{\partial_{q}}{\partial_{q}x_{\mu}}\mathrm{e}_{\mu}\frac{\partial_{q}}{\partial_{q}x_{\nu}},$ (58)
and substituting (56) and (57) into (58) we get
$D^{2}_{q}$ $=\ -2\delta_{\mu\alpha}\frac{\partial^{2}_{q}}{\partial_{q}x_{\mu}\partial_{q}x_{\alpha}}-\frac{\partial^{2}_{q}}{\partial_{q}x^{2}_{\nu}}+\delta_{\mu\alpha}\frac{\partial^{2}_{q}}{\partial_{q}x_{\mu}\partial_{q}x_{\alpha}},$ (59)
and for $\mu=\alpha$ finally we obtain $-\frac{\partial^{2}_{q}}{\partial_{q}x^{2}_{\mu}}-\frac{\partial^{2}_{q}}{\partial_{q}x^{2}_{\nu}}$, which completes the proof. ∎
According to above, we can said that this operator acts over a $q$-spinor function on the form $\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu},x_{\nu}))$. Therefore (55) can be expressed as
$D_{q}\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu},x_{\nu}))=\mathrm{e}_{\nu}\frac{\partial_{q}\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu},x_{\nu}))}{\partial_{q}x_{\mu}}+\mathrm{e}_{\mu}\frac{\partial_{q}\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu},x_{\nu}))}{\partial_{q}x_{\nu}}.$ (60)
and the derivatives $\frac{\partial_{q}\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu},x_{\nu}))}{\partial_{q}x_{\mu}}$ and $\frac{\partial_{q}\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu},x_{\nu}))}{\partial_{q}x_{\nu}}$ can be determined using the $q$-spinor chain rule defined by (51). We will consider the following example
@@ 2.2 The new $q$-differential operator for $q$-spinor variables
Let $\Psi(x_{\mu},x_{\nu})=\exp{ix_{\mu}}$ ) and $u^{1}_{\dot{2}}=q^{2}x_{\mu}$. Now, to obtain the differential operator (55), first we write $\Psi(u^{1}_{\dot{2}})$ as $\exp{iu^{1}_{\dot{2}}}$ ). Later we apply (51)
$\frac{\partial_{q}\Psi}{\partial_{q}x_{\mu}}$ $=\ \frac{\partial_{q}}{\partial_{q}u^{1}_{\dot{2}}}(\exp{iu^{1}_{\dot{2}}})\frac{\partial_{q}}{\partial_{q}x_{\mu}}(q^{2}x_{\mu}),$
[…]
$=\ q^{2}\frac{\exp{iqu^{1}_{\dot{2}}}-q\exp{iqu^{1}_{\dot{2}}}}{[qu]^{1}_{\dot{2}}-q(u^{1}_{\dot{2}})},$
finally we apply (55) to obtain
$D_{q}\Psi(u^{1}_{\dot{2}})=\mathrm{e}_{\nu}q^{2}\frac{\exp{iqu^{1}_{\dot{2}}}-q\exp{iqu^{1}_{\dot{2}}}}{[qu]^{1}_{\dot{2}}-q(u^{1}_{\dot{2}})}.$
The expression (55) can be applied to functions that not depends of $u^{\alpha}_{\dot{\beta}}$, That is, it depends only on the variables $x_{\mu},x_{\nu}$. To obtain this result, it is necessary to define the $q$- derivatives in terms of the $q$-spinor variable $x^{\alpha}$ and $x_{\dot{\beta}}$, and the elements $\mathrm{e}_{\mu},\mathrm{e}_{\nu}$ in the following definition
@@
Let us consider the function $\psi(x_{\mu},x_{\nu})$. The expressions:
@@
are the $q$-derivatives respect to $x_{\mu}$ and $x_{\nu}$ in terms of $q$- spinor variables $x_{\dot{\beta}},x^{\alpha}$ and the elements $\mathrm{e}_{\mu},\mathrm{e}_{\nu}$.
Let us consider the following remark.
@@
Taking into account the above claim, we can said that the differential operator not apply only over the functions $\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu}))$, but also over the function $\psi(x_{\mu},x_{\nu})$, and therefore the operator (55) also depends of the variables $x_{\mu}$ and $x_{\nu}$.
@@
Consider the following function $\psi(x_{\mu},x_{\nu})=qx_{\nu}x_{\dot{\beta}},\dot{\beta}=\dot{2}$. Now, to determine the derivatives, we apply (61) and (62)
$\frac{\partial_{q}\psi}{\partial_{q}x_{\mu}}$ $=0$
$\frac{\partial_{q}\psi}{\partial_{q}x_{\nu}}$ $=\ \frac{q(x_{\nu}+q\mathrm{e}_{\nu}x_{\dot{2}})x_{\dot{2}}-qx_{\nu}x_{\dot{2}}}{x_{\dot{2}}}$
$=\ q^{2}\mathrm{e}_{\nu}x_{\dot{2}},$
thus the operator (55) is expressed as
@@ 2.3 The $q$-Dirac differential operator
The $q$-analogue of (42) can be expressed as
@@
We can now formulate our main results in the following propositions.
@@ 2.4 The $q$-differential operators for $q$-spinor variables
Let $\Psi$ be a function on the $q$-spinor variables. The operator (55) for $q$-spinor variables $D^{q}\Psi$ is given by
@@
Multiplying the left-hand side by $\mathrm{e}_{\nu}$ in (54), we have
@@
adding (66) with (67) and considering (55), finally results
@@
The above proof implies the following relation
@@
If $\Psi$ also depends of $x_{\nu}$, then
@@
consequently the $q$-differential for $u^{\alpha}_{\dot{\beta}}$ is defined as
@@
Let $\Psi$ be a function on the $q$-spinor variables. The Dirac operator for $q$-spinor variables $D_{\mu}^{q}\Psi$ is given by
@@
Multiplying the left-hand side by $\gamma_{\mu}$ in (54), we have
@@
The above proof implies the following relation
@@
consequently the Dirac $q$-differential for $u^{\alpha}_{\dot{\beta}}$ is defined as
@@
Now, by (52), (64) and (72), we have the follows $q$-spinor integral formulas showed in the following subsection.

### v2

@@ 2 $q$-Differential Operators for $q$-Spinor Variables
The aim of this section is to define a $q$-differential operator for $q$-spinor variables. To begin with, we introduce a fundamental rule of the $q$-spinor differential calculus, namely the $q$-spinor chain rule.
@@ 2.1 The $q$-Spinor Chain Rule
Let $\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu}))$ be a $q$-spinor function. Then the $q$-spinor chain rule is given by
$\frac{\partial^{q}\Psi}{\partial^{q}x_{\mu}}=\frac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}\cdot\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial_{q}x_{\mu}}.$ (1)
@@
Consider the $q$-derivative of the composite function:
$\frac{\partial^{q}\Psi}{\partial^{q}x_{\mu}}=\frac{\Psi\left((qu)^{\alpha}_{\dot{\beta}}(x_{\mu})\right)-q\,\Psi\left(u^{\alpha}_{\dot{\beta}}(x_{\mu})\right)}{qx_{\mu}-qx_{\mu}}.$ (2)
Multiplying and dividing by the non-vanishing quantity $(qu)^{\alpha}_{\dot{\beta}}(x_{\mu})-q\,u^{\alpha}_{\dot{\beta}}(x_{\mu})$, we may rewrite (2) as
$\frac{\partial^{q}\Psi}{\partial^{q}x_{\mu}}=\frac{\Psi\left((qu)^{\alpha}_{\dot{\beta}}(x_{\mu})\right)-q\,\Psi\left(u^{\alpha}_{\dot{\beta}}(x_{\mu})\right)}{(qu)^{\alpha}_{\dot{\beta}}(x_{\mu})-q\,u^{\alpha}_{\dot{\beta}}(x_{\mu})}\cdot\frac{(qu)^{\alpha}_{\dot{\beta}}(x_{\mu})-q\,u^{\alpha}_{\dot{\beta}}(x_{\mu})}{qx_{\mu}-qx_{\mu}}.$ (3)
Denoting the first factor as $\frac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}$ and the second as $\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial^{q}x_{\mu}}$, we obtain
$\frac{\partial^{q}\Psi}{\partial^{q}x_{\mu}}=\frac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}\cdot\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial^{q}x_{\mu}},$ (4)
which establishes the desired result. ∎
We now proceed to define a novel $q$-differential operator acting on $q$-spinor variables over an orthonormal basis of $\mathds{R}^{n}$. This operator differs from the classical Dirac and Cauchy–Riemann operators considered in [3, 4, 6, 7, 9].
2.2 The new $q$-differential operator for $q$-spinor variables
The motivation arises from the construction of a differential operator satisfying
$D_{q}^{2}=-\frac{\partial^{2}_{q}}{\partial_{q}x^{2}_{\mu}}-\frac{\partial^{2}_{q}}{\partial_{q}x^{2}_{\nu}},$
for all $\mu,\nu=1,2,\dotsc,n$, over an orthonormal basis of $\mathds{R}^{n}$.
Proposition 2.2.
Let $\left\{\mathrm{e}_{1},\mathrm{e}_{2},\dotsc,\mathrm{e}_{n}\right\}$ be an orthonormal basis of $\mathds{R}^{n}$. The $q$-differential operator $D^{q}$ defined by
satisfies the relations:
$\mathrm{e}_{\mu}^{2}\frac{\partial_{q}^{2}}{\partial_{q}x_{\nu}^{2}}+\mathrm{e}_{\nu}^{2}\frac{\partial_{q}^{2}}{\partial_{q}x_{\mu}^{2}}$ $=-2\delta_{\mu\alpha}\frac{\partial_{q}^{2}}{\partial_{q}x_{\mu}\partial_{q}x_{\alpha}}-\frac{\partial_{q}^{2}}{\partial_{q}x_{\nu}^{2}},$ (6)
$\mathrm{e}_{\mu}\frac{\partial_{q}}{\partial_{q}x_{\nu}}\mathrm{e}_{\nu}\frac{\partial_{q}}{\partial_{q}x_{\mu}}+\mathrm{e}_{\nu}\frac{\partial_{q}}{\partial_{q}x_{\mu}}\mathrm{e}_{\mu}\frac{\partial_{q}}{\partial_{q}x_{\nu}}$ $=\delta_{\mu\alpha}\frac{\partial_{q}^{2}}{\partial_{q}x_{\mu}\partial_{q}x_{\alpha}},\quad\mu,\nu,\alpha=1,2,\dotsc,n.$ (7)
Proof.
The proof relies on the observation that the square of (5) is equivalent to $-\frac{\partial_{q}^{2}}{\partial_{q}x_{\mu}^{2}}-\frac{\partial_{q}^{2}}{\partial_{q}x_{\nu}^{2}}$. Computing $D_{q}^{2}$ gives
$D_{q}^{2}$ $=\mathrm{e}_{\mu}^{2}\frac{\partial_{q}^{2}}{\partial_{q}x_{\nu}^{2}}+\mathrm{e}_{\nu}^{2}\frac{\partial_{q}^{2}}{\partial_{q}x_{\mu}^{2}}+\mathrm{e}_{\mu}\frac{\partial_{q}}{\partial_{q}x_{\nu}}\mathrm{e}_{\nu}\frac{\partial_{q}}{\partial_{q}x_{\mu}}+\mathrm{e}_{\nu}\frac{\partial_{q}}{\partial_{q}x_{\mu}}\mathrm{e}_{\mu}\frac{\partial_{q}}{\partial_{q}x_{\nu}}.$ (8)
Substituting (6) and (7) into (8) yields
$D_{q}^{2}$ $=-2\delta_{\mu\alpha}\frac{\partial_{q}^{2}}{\partial_{q}x_{\mu}\partial_{q}x_{\alpha}}-\frac{\partial_{q}^{2}}{\partial_{q}x_{\nu}^{2}}+\delta_{\mu\alpha}\frac{\partial_{q}^{2}}{\partial_{q}x_{\mu}\partial_{q}x_{\alpha}}.$ (9)
For $\mu=\alpha$, we obtain the claimed identity
$D_{q}^{2}=-\frac{\partial_{q}^{2}}{\partial_{q}x_{\mu}^{2}}-\frac{\partial_{q}^{2}}{\partial_{q}x_{\nu}^{2}},$
@@ 2.2 The new $q$-differential operator for $q$-spinor variables
According to the above result, this operator acts on $q$-spinor functions of the form $\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu},x_{\nu}))$. Therefore, the operator (5) can be rewritten as
$D_{q}\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu},x_{\nu}))=\mathrm{e}_{\nu}\frac{\partial_{q}\Psi(u^{\alpha}_{\dot{\beta}})}{\partial_{q}x_{\mu}}+\mathrm{e}_{\mu}\frac{\partial_{q}\Psi(u^{\alpha}_{\dot{\beta}})}{\partial_{q}x_{\nu}},$ (10)
where the $q$-derivatives are computed using the chain rule (1) for $q$-spinor variables.
@@
Let $\Psi(x_{\mu},x_{\nu})=\exp(ix_{\mu})$ and define $u^{1}_{\dot{2}}=q^{2}x_{\mu}$. Then, $\Psi(u^{1}_{\dot{2}})=\exp(iu^{1}_{\dot{2}})$. Using (1), we compute:
$\frac{\partial_{q}\Psi}{\partial_{q}x_{\mu}}$ $=\frac{\partial_{q}}{\partial_{q}u^{1}_{\dot{2}}}(\exp(iu^{1}_{\dot{2}}))\cdot\frac{\partial_{q}}{\partial_{q}x_{\mu}}(q^{2}x_{\mu})$
$=q^{2}\cdot\frac{\exp(iqu^{1}_{\dot{2}})-q\exp(iu^{1}_{\dot{2}})}{(qu)^{1}_{\dot{2}}-qu^{1}_{\dot{2}}}\cdot\frac{\partial_{q}x_{\mu}}{\partial_{q}x_{\mu}}$
$=q^{2}\cdot\frac{\exp(iqu^{1}_{\dot{2}})-q\exp(iqu^{1}_{\dot{2}})}{(qu)^{1}_{\dot{2}}-qu^{1}_{\dot{2}}},$
and applying (5), we obtain:
$D_{q}\Psi(u^{1}_{\dot{2}})=\mathrm{e}_{\nu}q^{2}\cdot\frac{\exp(iqu^{1}_{\dot{2}})-q\exp(iqu^{1}_{\dot{2}})}{(qu)^{1}_{\dot{2}}-qu^{1}_{\dot{2}}}.$
The expression (5) also applies to functions that depend directly on the variables $x_{\mu}$ and $x_{\nu}$, without requiring the spinor composition. For this case, we define the $q$-derivatives in terms of the $q$-spinor variables $x^{\alpha}$ and $x_{\dot{\beta}}$ as follows:
@@
Let $\psi(x_{\mu},x_{\nu})$ be a scalar function. The $q$-derivatives with respect to $x_{\mu}$ and $x_{\nu}$ are defined by
@@
where $x^{\alpha}$ and $x_{\dot{\beta}}$ are $q$-spinor variables, and $\mathrm{e}_{\mu},\mathrm{e}_{\nu}$ belong to the orthonormal basis of $\mathds{R}^{n}$.
@@
In view of the above, we conclude that the operator (5) not only acts on spinor-composed functions $\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu}))$, but also on scalar functions $\psi(x_{\mu},x_{\nu})$. Hence, it inherently depends on the variables $x_{\mu}$ and $x_{\nu}$.
@@
Consider the scalar function $\psi(x_{\mu},x_{\nu})=qx_{\nu}x_{\dot{\beta}}$, with $\dot{\beta}=\dot{2}$. Applying (11) and (12), we obtain:
$\frac{\partial_{q}\psi}{\partial_{q}x_{\mu}}$ $=0,$
$\frac{\partial_{q}\psi}{\partial_{q}x_{\nu}}$ $=\frac{q(x_{\nu}+q\,\mathrm{e}_{\nu}x_{\dot{2}})x_{\dot{2}}-qx_{\nu}x_{\dot{2}}}{x_{\dot{2}}}=q^{2}\mathrm{e}_{\nu}x_{\dot{2}}.$
Thus, the operator (5) yields:
@@ 2.3 The $q$-Dirac differential operator
The $q$-analogue of the Dirac operator is defined by
@@
We are now ready to state our main results in the following propositions.
@@ 2.4 The $q$-differential operators for $q$-spinor variables
Let $\Psi$ be a function of the $q$-spinor variables. The operator (5) for $q$-spinor variables $D^{q}\Psi$ is given by
@@
Multiplying the left-hand side by $\mathrm{e}_{\nu}$ in (4), we obtain
@@
Adding (16) and (17) and considering (5), we finally obtain
@@
The above proof implies the following relation:
@@
If $\Psi$ also depends on $x_{\nu}$, then
@@
consequently, the $q$-differential for $u^{\alpha}_{\dot{\beta}}$ is defined as
@@
Let $\Psi$ be a function of the $q$-spinor variables. The Dirac operator for $q$-spinor variables $D_{\mu}^{q}\Psi$ is given by
@@
Multiplying the left-hand side by $\gamma_{\mu}$ in (4), we obtain
@@
The above proof implies the following relation:
@@
consequently, the Dirac $q$-differential for $u^{\alpha}_{\dot{\beta}}$ is defined as